-[3+4]+[-5]+[6-10]=
Answer by midwood_trail(310)

-[3+4]+[-5]+[6-10]=
-|7| + 5 + |-4| =
-7 + 5 + 4 =
-7 + 9 = 2
The absolute value is distance and so it is always positive.